Two straight line OX and OY cut at right angles and from P.a point with in the angle XOY,perpendiculars PM,PN are drawn to OX and OY respectively. Plot the locus of P when 1)PM+PN is constant (6cm)2)PM-PN is constant (3cm)

Grade:12th pass

1 Answers

Sai
5 years ago
If PM+PN=6 Then the points satisfying the given condition are(1,5),(2,4),.....etc.slome m=(4-5)÷(2-1)=-1.The eq`n is y-5=-1(x-1)x+y=6||ly if PM-PN=3x-y+3=0